BI-Linac4 meeting of Wednesday 25th October 2006 Minutes by M. Pasini and K. Hanke

Present: M. Pasini, K. Hanke, U. Raich, L. Soby, C. Dutriat, E. Bravin, G. Tranquille F. Lenardon

Agenda: 1. Follow-up of the previous meeting. 2. Finalize design, time-lines and cost estimate for the various devices in Linac4. [all equipment specialists] 3. LEBT beam current transformer status [Lenardon] and general planning revision 4. Chopper line, slow wire scanners refurbishing [Raich] 5. Integration of different component in the Linac. [all specialists] ***

1. M. Pasini reported that concerning the cost estimate table he had received only the contribution of L.Soby. U.Raich reported that he has just sent the old estimation and that the new table should be filled starting from those data. E.Bravin suggested that the section regarding the emittance scanner should be split in 4 parts since each device is quite different from each other. M.Pasini will gather all the contribution and will put them into the Excel table. Afterwards he will distribute the table for internal checking.

As for the beam test of the BSHM device in Orsay, K.Hanke reported that it is planned for the second half of next year. E.Bravin could be available for following the test as well.

2. All specialists underlined the fact of the lack of time to look in more in detail into the design of the several instruments. The dimension that they can provide is only at the level of order of magnitude, with some exception. L.Soby has provided the 2 designs of the Pus. The problem of the alignment and of the mechanical support was brought up. The answer to this problem is not possible since a design of the instruments does not exist. L.Soby asked about the possibility of having an external RF reference. M.Pasini answered that this is not a problem. M.Pasini asked if the old instruments proposed by the BI group are compatible with the Linac4 beam. E.Bravin and U.Raich answered that there should be no problem.

3. As for the LEBT beam current transformer F.Lenardon has ordered the tore and done 2 designs of the ceramics. Two spares for the ceramics will be ordered. The work is progressing.

No progress on the Faraday cups.

As for the LEBT planning K. Hanke reported that the old planning presented by R.Scrivens is not feasible and a new planning has to be done in agreement with the general planning kept up to date by C.Rossi. K.Hanke volunteered to meet C.Rossi to discuss about the new plan and to report to the meeting. action: K.Hanke.

4. K.Hanke reported that for the chopper line two slow wire scanner have been found a position in the line been identified. They need some refurbishing of the electronics and the software. U.Raich will look into this with J.J.Gras. action: U.Raich, J.J.Gras.

5. M. Pasini reported that Y. Cuvet is working on the general plan of Linac4. This will give the opportunity to have a look of the global aspect of the machine and to check about the number of the components needed in the linac.

-- KlausHanke - 13 Sep 2007

Edit | Attach | Watch | Print version | History: r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r2 - 2008-06-30 - FrankGerigk
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    SPL All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright